# Non-Gaussianity in SMICA

### v1

@@ 5.1 The multivariate Edgeworth expansion
In order to incorporate this extra information in the SMICA model, which is a component separation method that works in the harmonic space, it is a natural step to relax the assumption of Gaussianity to include higher order $n$-point correlators. We make use of the Multivariate Edgeworth Expansion (MEE) [26, 27, 28, 29, 30]: a generic PDF $P(a_{lm}^{d})$, describing the distribution of the spherical harmonic coefficients of the sky in temperature T, E-mode or B-mode polarization independently, at a given frequency indicated by the index $d$, can be expanded around a multivariate gaussian $G(a^{d}_{lm}|\textbf{C})$ as
@@ 6.4 E-mode polarizations results
In order to compare the results of our fit we have also performed the "standard" non-gaussian analysis reconstructing via the Generalised Least Square (GLS) filter the CMB and noise maps after component separation and applying the usual binned bispectrum estimator on them. Under gaussian statistics and without other prior assumptions on the astrophysical components, given the linear multi-frequency and multi-component model in eq. (3.1), one can use a generalized least-square solution in harmonic space [31, 1]
@@
In figure 10 we show the results obtained with our multi-frequency multi-component binned bispectrum estimator and with the standard binned bispectrum estimator for a masked sky using the same mask of section 4. In the two cases, we recover consistent results: $f_{\text{NL}}=99\pm 290$ for the first one and $f_{\text{NL}}=6\pm 280$ for the second. These large error bars might surprise since the current constraint on $f_{\text{NL}}$ for the local shape is $\sigma(f_{\text{NL}})=5$ given by the analysis of the Planck data [17], but it is important to note that that limit was obtained via a combined temperature and E-mode analysis going up to $\ell_{\text{max}}=2500$, while here we have E only and $\ell_{\text{max}}=100$.
As in section 6.3 we show the results obtained for the squeezed configuration in figure 11 and figure 12 and the results for the equilateral configuration in figure 13 and figure 14. In the E-mode polarization case the significance of the squeezed configuration over the equilateral one is much more evident for the two foregrounds. Indeed, the squeezed configuration is recovered without any bias for both dust and synchrotron, while the equilateral bispectrum of the foregrounds has a lower amplitude that tends to zero.
@@
Figure 10: Comparison of the results for $f_{\text{NL}}$ between the standard binned bispectrum estimator and the multi-frequency multi-component binned bispectrum estimator for the 400 simulations in the masked-sky case.
@@ 7 Conclusions
Based on these results, we developed an independent bispectrum estimator: since a combined power spectrum and bispectrum analysis has shown no improvement for the power spectrum estimation, we developed a multi-frequency likelihood for the binned bispectrum and performed a conditioned estimation of the bispectrum with the power spectrum obtained independently by SMICA. With this estimator, we were able to recover the correct bispectra of the two foregrounds included in our simulations, dust and synchrotron, while also recovering coherent constraints on local primordial non-Gaussianity with respect to the standard post component separation analysis. However, our approach allows for a better handling of the foregrounds as it applies an optimal combination of the channels.

### v2

@@ 5.1 The multivariate Edgeworth expansion
In order to incorporate this extra information in the SMICA model, which is a component separation method that works in the harmonic space, it is a natural step to relax the assumption of Gaussianity to include higher order $n$-point correlators. We make use of the Multivariate Edgeworth Expansion (MEE) [26, 27, 28, 29, 30]: a generic probability density function (PDF) $P(a_{lm}^{d})$, describing the distribution of the spherical harmonic coefficients of the sky in temperature T, E-mode or B-mode polarization independently, at a given frequency indicated by the index $d$, can be expanded around a multivariate gaussian $G(a^{d}_{lm}|\textbf{C})$ as
@@ 6.4 E-mode polarizations results
In order to compare the results of our fit we have also performed the "standard" non-gaussian analysis, reconstructing via the Generalised Least Square (GLS) filter the CMB and noise maps after component separation and applying the usual binned bispectrum estimator on them. Under gaussian statistics and without other prior assumptions on the astrophysical components, given the linear multi-frequency and multi-component model in eq. (3.1), one can use a generalized least-square solution in harmonic space [31, 1]
@@
In figure 10 we show the results obtained with our multi-frequency multi-component binned bispectrum estimator and with the standard binned bispectrum estimator for different masks. In the two cases, we recover consistent results. This was expected as we are working with a simplified foreground model, in which case the GLS filter efficiently removes foreground residuals from the CMB map. With a more complex sky, our approach is expected to be more robust and more efficient in detecting deviations from the idealized model, as the foregrounds are already included in the analysis. The large error bars in figure 10 might surprise since the current constraint on $f_{\text{NL}}$ for the local shape is $\sigma(f_{\text{NL}})=5$ given by the analysis of the Planck data [17], but it is important to note that that limit was obtained via a combined temperature and E-mode analysis going up to $\ell_{\text{max}}=2500$, while here we have E only and $\ell_{\text{max}}=100$. Moreover, in this analysis we have not included the linear correction term, which is required to obtain optimal error bars if isotropy is broken, as is the case here because of the mask and anisotropic foregrounds.
As in section 6.3 we show the results obtained for the foregrounds in the squeezed configuration in figure 11 and figure 12 and the results in the equilateral configuration in figure 13 and figure 14. In the E-mode polarization case the significance of the squeezed configuration over the equilateral one is much more evident for the two foregrounds. Indeed, the squeezed configuration is recovered without any bias for both dust and synchrotron, while the equilateral bispectrum of the foregrounds has a lower amplitude that tends to zero.
@@
Figure 10: Comparison of the results for $f_{\text{NL}}$ and their error bars between the standard binned bispectrum estimator and the multi-frequency multi-component binned bispectrum estimator for the 400 simulations for masks with a unmasked sky fraction of 60%, 70%, 80% and 90%.
@@ 7 Conclusions
Based on these results, we developed an independent bispectrum estimator: since a combined power spectrum and bispectrum analysis has shown no improvement for the power spectrum estimation, we developed a multi-frequency likelihood for the binned bispectrum and performed a conditioned estimation of the bispectrum with the spectral parameters obtained independently by SMICA. With this estimator, we were able to recover the correct bispectra of the two foregrounds included in our simulations, dust and synchrotron, while also recovering coherent constraints on local primordial non-Gaussianity with respect to the standard post-component separation analysis. However, our approach allows for a better handling of the foregrounds, in particular in the case of more complicated foreground models where we expect our method to outperform the standard approach.